OpenStudy (anonymous):

solve for y 2xsquared+2ysquared=5

OpenStudy (anonymous):

\[2x^2+2y^2=5\] \[2y^2=5-2x^2\] \[y^2=\frac{5}{2}-\frac{2x^2}{2}\] \[y=\pm \sqrt{\frac{5}{2}-x^2}\]

OpenStudy (anonymous):

thanks! can you do 3x squareroot2(3-5y+sinx=3xy+8 to find y